It's been a long time since I've done Algebra and I don't know how to help my daughter with this one....
The three most populous countries of the world are China, India, and the Soviet Union, with a combined population of 1,650 million. Two hundred million more people live in China than in India, and the combined population of India and the Soviet Union is 50 million more than that of China. How many people live in each country.
I'd like to see how this is written out in equation form....not just get the answer.

Answer by stanbon(75887)

The three most populous countries of the world are China, India, and the Soviet Union, with a combined population of 1,650 million. Two hundred million more people live in China than in India, and the combined population of India and the Soviet Union is 50 million more than that of China. How many people live in each country.
-------------------------------------
Let population of China be "c", of india be "i", of Russia be "r".
EQUATIONS:
c + i + r = 1650 million
c - i = +200
i+r = c + 50
----------------
Rearrange the equations:
1c + 1i + ir = 1650
1c - 1i + 0r = 200
1c - 1i - 1r = -50
-------------------------
Using a TI calculator I get:
c = 800 million
i = 600 million
r = 250 million
===================
Cheers,
Stan H.